# Anisotropic anomalous diffusion in microgravity dusty plasma. Part One: Nonequilibrium Statistical Analysis

### v2

@@ 4.2 Fitting Techniques
Here $A,v_{th},v_{0}$, and $q$ are the fitting parameters for the velocity histograms. The fits to the position histograms used the same distributions where $v_{th}$ was replaced by $D$ and $(v_{r}-v_{r0})^{2}$ was replaced by $(r(\tau)-r_{0})^{2}$. To minimize fitting errors, the normalization in Eq. 6 was set to $1$. Dropping this normalization has a negligible effect. For all cases, it was found that the axial component of the velocities (displacements) is best approximated by a single q-Gaussian, while the radial component was best approximated by a Bi-q-Gaussian. The value of $q$ found from the position and velocity histogram fits are denoted as $q_{p}$ and $q_{v}$, respectively. Data obtained from particle displacements/velocities along the axial direction is labeled by a subscript $\parallel$, while cross-field direction data have a subscript $\perp$.
@@ 5.1 Fits to MSD Plots
To better understand the role of directional anisotropy, we calculated MSDs from particle displacements along $\parallel$ and across $\perp$ the direction of the external electric field. Separate fits were performed at different time delays to assess the role of distinct physical processes. Only data corresponding to time delays smaller than $10s$ was used for the fits. At larger time delays, the standard deviation from the mean MSD increases considerably due to decreasing number of data points. Figure 5 shows representative plots of $MSD_{\parallel}$ and $MSD_{\perp}$ for the 70Pa, 1mA case. For all pressure-current conditions the $MSD_{\parallel}$ has a magnitude much greater than $MSD_{\perp}$ and the plots of $MSD_{\parallel}$ look almost identical to those of the combined MSDs in Fig. 5. The $MSD_{\parallel}$ for all conditions has a positive concavity at all time delays, suggesting superdiffusion. The one exception is the (30 Pa, 0.35 mA) case, which has a brief negative concavity at its start. This dataset had the most noisy trajectories and the least number of data points at long time delays. Thus, the deviation in observed behavior may be due to the poorer quality of the data. The $MSD_{\perp}$ for all cases exhibits a brief negative concavity for a delay period of about two seconds, followed by a positive concavity at lager time scales. This suggests that there may be a trapping mechanism causing sub-diffusive behavior at time scales smaller than $2s$. These effects are showcased in the plots in Fig 6.
@@
Figure 7: $\alpha_{\parallel}$ and $\alpha_{\perp}$ for all pressure-current cases.
The exponent $\alpha$ extracted from $MSD_{\parallel}$ and $MSD_{\perp}$ at time delay between $4~{}s$ and $5~{}s$ is plotted for all cases in Fig. 7. This time period was chosen as it has the most consistent slope fitting while not being too short of a time delay. Recall that Tsallis’ theory claims that the diffusion is a Lévy process when $q>5/3$. Using the scaling relation $MSD=\langle r^{2}\rangle\propto\tau^{\frac{2}{3-q_{p}}}$ [63] yields the criterion $\alpha>3/2$ for a Lévy process. Note that this scaling is valid only for $q_{p}$ obtained from the displacement distributions. A dashed line on Fig. 7 marks the location of $\alpha=3/2$, indicating that almost all distributions for the parallel cases are a Lévy processes, while the perpendicular direction are not. Later we compare this value of $\alpha$ to that found from $q_{p}$.
@@ 5.2 Fits to Displacement Histograms
Representative histograms of the displacements are shown in Fig. 9 for 70 Pa 0.7 mA case. The histograms of displacements parallel to the electric field are best approximated by a single q-Gaussian distribution, while the histograms of cross-field displacements are best approximated by a Bi-q-Gaussian distribution in most cases. These are displacements of all particles at time delay $\tau=5$. The directional difference of the distributions type implies there is anisotropic diffusion. The nonextensive $q_{p}$ parameters extracted from the displacement histograms are plotted in Figure 8.
Figure 8: Position distribution fit example from case 70 Pa 0.7 mA for displacements in a) parallel direction and b) cross field direction plotted in logarithmic scale. The data for the position distributions is the positive squared value of $(r(\tau)-r_{0})^{2}$ instead of the velocity distributions $v_{r}$. Since the data of $(r(\tau)-r_{0})^{2}$ is already squared we fit it to a q-exponential instead of a q-Gaussian.
@@
Notice that the coefficients for the parallel direction are again larger than for the perpendicular direction. Dashed line indicates a Lévy process. The coefficients extracted from the histogram of parallel displacements suggest that two of the 70Pa cases are consistent with a Lévy process. A crossover to a Lévy process is also seen for the cross-field direction in the 70 Pa, 1mA case if a single q-Gaussian is fitted to the histogram of cross-field displacements. However, the more accurate Bi-q-Gaussian fit to the cross-field displacement histograms suggest that the process is likely superdiffusive but not Lévy. These results are in qualitative agreement with the conclusions drown from the $\alpha$ coefficients extracted from the MSD fits at time delay $\tau=5$ shown in Figure 7. It seems that higher pressures create larger spread in the coefficients and exhibit a larger difference between the parallel and perpendicular directions suggesting a more pronounced anisotropic effect. The values for the diffusion $D\left(\frac{\mu m^{2}}{s}\right)$ calculated from Eq. 10, are shown in Fig. 10. Here we can see that, in the direction parallel to the field, an increase in pressure causes diffusion to increase. However, in the cross-field direction, the interpretation is less straightforward. $D_{\perp 1}$ and $D_{\perp}$ generally decrease with pressure with some exceptions. $D_{\perp 2}$ does not show a clear trend with pressure or current.
@@
Next, we examined how the exponent $\alpha$ found directly from MSD fits compares to $\alpha$ calculated from the displacement histogram fits using the scaling in Eq. 7. Figure 11 shows the percent difference $\frac{\alpha_{MSD}-\alpha_{q_{p}}}{\alpha_{q_{p}}}$. We can see that on average the error is lower, indicated by the dashed line at 20%, for the cases where a Bi-q-Gaussian fit was used rather than a single q-Gaussian. This seems to be in agreement with findings from [30, 14].
@@ 5.3 Velocity Distribution Plots
Next we examine Fig. 12, which shows representative histograms and distribution fits for the velocity components parallel and perpendicular to the direction of the external electric field. These plots show that a Gaussian or a Maxwellian distribution (black line) does not approximate the ‘fat’ tails of the histograms, while a q-Gaussian (magenta line) and a Bi-q-Gaussian (green line) provide good fits to the $\parallel$ and the $\perp$ velocity histograms, respectively. It is again observed that the distribution features change substantially with direction, suggesting anisotropic dust diffusion caused by the external electric field. Similar to the displacement data, we observe that the velocity histograms in the cross-field direction are best approximated by a Bi-q-Gaussian fit, suggesting a superposition of two distinct diffusion processes. These trends in histogram shapes are observed for all pressure-current cases. This will be further discussed in Sec. 6.2.
@@
Kinetic temperature is typically defined as the variance of a Maxwellian distribution. Since the velocity histograms for the dust in the PK-4 experiment are best approximated by a non-Maxwellian, the notion of "temperature" here is not well defined. The range of kinetic temperatures found from the Maxwellian fits (Fig. 14) is $2000K-6500K$, or about $0.2eV-0.6eV$, suggesting that the majority of the dust particles have very high kinetic energy even though the experiment is at room temperature and the pressure is low. Anomalously high dust temperature is a known phenomena that is caused by electrostatic fluctuations [69]. In Fig. 15, we used Eq. 10, $T_{q}(\frac{5q-3}{2})=T_{Mq}$ (from [56]), since it provides a relation between the temperature or thermal velocity found from a q-Gaussian and to a ‘Maxwellian-like’ temperature.
@@
As shown in Figure 12 b), the perpendicular velocity histograms are best approximated by a Bi-q-Gaussian fit, where the the bulk distribution (the one with the larger variance and smaller $q_{v}$ value) more closely resembles a Gaussian and, thus, can be used to extract the thermal velocity. The second distribution in the Bi-q-Gaussian accounts for the "fat tails". Thus, in Fig. 15, $T_{\perp 1}$ is extracted from a Gaussian-like sub-population distribution, while $T_{\perp 2}$ is obtained from fitting to a ’tail-halo’ distribution. It may seem unintuitive why the Gaussian sub-population would have higher temperature values, while the ’halo-tails’ have lower temperature values. This will be discussed in 6.2. Essentially we convert a variance from a distribution which has some ’tailed-ness’ to a variance which does not have any, so that we can use the same definition of temperature. We argue that this is not a different treatment of temperature in dusty plasma, but an attempt at calculating temperature more carefully with nonextensive statistics. This allows us to derive more accurate dust temperatures for the PK-4 dusty plasma with values in the range $2500K-5500K$, or about $0.2eV-0.5eV$, which is smaller than the typically calculated values of ${}10-300eV$ [69].
@@ 5.4 Domain Velocity Distribution Plots
Here we provide physical arguments why the Bi-q-Gaussian fit for the perpendicular histograms is needed. First, we consider the domain separation on Fig. 3 A) which was made for the NHDS method and reconstruct the velocity histogram in each domain. Figure 16 shows the velocity distribution for domain 11 (dark blue in Fig. 3 A) in both the parallel and perpendicular directions for from the 30 Pa 0.7 mA case.
@@
3: The three tables show $q_{v}$ placed in the table at the corresponding domain showing a spatial map of the equilibrium.
We use the newly-founded q-Gaussian fits to calculate temperature in each domain using Eq. 10. Figure 17 a) $T_{\|}$ and b) $T_{\perp}$ show temperatures in units of Kelvin for 20 domains calculated for the 30 Pa 0.7 mA case. Each domain contained about 500 dust tracks, a minimum of 10 time frames per trajectory, which yields statistically significant amount of data for the velocity fits. The data indicates that temperatures gradients exist in the PK4 system. As seen from Eq. 13, the temperature gradient and and the gradient in potential are related through the $q_{v}$ parameter. MD simulations of ions and dust in the PK-4, with conditions closely resembling the 70 Pa 0.7 mA case, reveal gradients in the electric potential surrounding each dust particles, as shown in Fig. 18 [38]. Examination of Figures 17 and 18 suggests that the dust clouds in PK-4 exhibit both temperature and electric potential gradients, in addition to q-Gaussian velocity distribution functions. This provides a qualitative argument for applying Eq. 12, which is dependent on the ratio of the thermophoretic force over the electric field force. Providing a quantitative validation for Eq. 13 is beyond the scope of the present paper and will be explored in future work.
Figure 17: $T_{\|}$ and $T_{\perp}$ calculated in each domain showing temperature gradients (30 Pa 0.7 mA).
@@ 6.1 MSD and Displacement Histograms
When analyzing the MSD plots fitting (Fig. 7), it seems like both $\alpha_{\parallel}$ and $\alpha_{\perp}$ increase with pressure for dc currents 0.35 mA and 0.7 mA. The trends are less obvious at 1 mA. These fits also indicate that none of the extracted $\alpha_{\perp}$ exponents correspond to a Lévy process, while all data for $\alpha_{\parallel}$ suggest a Lévy processes. Meanwhile, the $q_{p}$ coefficients extracted from the position histogram fits (Fig. 9) also suggest that $q_{p\|}$ increases with pressure (i.e., enhanced superdiffusion), in agreement with the $\alpha_{\parallel}$ trends. However, only the highest pressure-currents case (70 Pa 1mA) in Fig. 9 indicates a Lévy processes (i.e., $q>5/3$). The $q_{p\perp}$ values extracted from single q-Gaussian fits show similar trends for the same currents as $\alpha_{\perp}$ but the 70Pa 1mA case suggests a Lévy process, unlike $\alpha_{\perp}$. The Bi-q-Gaussian fits yield $q_{p\perp 2}$ values that exhibit closer similarity to the $\alpha_{\perp}$ trends. We can also see from Figure 11 that the fits for $q_{p\perp 2}$ and $q_{p\perp 1}$ yield lower percent error between the predicted and measured $\alpha_{\perp}$, strengthening the argument that the Bi-q-Gaussian is a more appropriate fit. The calculated diffusion constants (Fig. 10) show an overall increase in parallel diffusion with pressure, while the perpendicular diffusion seems to decrease with pressure when a single q-Gaussian fit is used. These trends are not as clear for $D_{\perp 1}$ and $D_{\perp 2}$ which were obtained from a Bi-q-Gaussian fit. Both the MSD fits and the displacement histogram fits indicate a pronounced difference in the diffusion regime when comparing the direction along the electric field versus the cross-field direction, which can be attributed to an anisotropy in the ion wake-mediated dust-dust interaction potential.

### v3

@@ 4.2 Fitting Techniques
Here $A,v_{th},v_{0}$, and $q$ are the fitting parameters for the velocity histograms. The fits to the position histograms used the same distributions where $v_{th}$ was replaced by $D$ and $(v_{r}-v_{r0})^{2}$ was replaced by $(r(\tau)-r_{0})^{2}$. To minimize fitting errors, the normalization in Eq. 7 was set to $1$ and the fitting function from Eq. 10 from [61] was used. Dropping this normalization has a negligible effect. For all cases, it was found that the axial component of the velocities (displacements) is best approximated by a single q-Gaussian, while the radial component was best approximated by a Bi-q-Gaussian. The value of $q$ found from the position and velocity histogram fits are denoted as $q_{p}$ and $q_{v}$, respectively. Data obtained from particle displacements/velocities along the axial direction is labeled by a subscript $\parallel$, while cross-field direction data have a subscript $\perp$. The kinetic temperature was calculated from the thermal velocity parameters, while it is defined as
[…]
This still requires the distribution curve $f(v)$ which we are not assuming to be Maxwellian but finding from the histogram. It is therefore faster to calculate temperature from the fitted variance which is the thermal velocity than to integrate the function to find the second moment of the distribution which is the variance or Temperature. Also, the integral does not converge for $q>5/3$ as discussed earlier, however for $q<5/3$ both methods were compared and result was a small difference of a few Kelvin.
@@ 5.1 Fits to MSD Plots
To better understand the role of directional anisotropy, we calculated MSDs from particle displacements along $\parallel$ and across $\perp$ the direction of the external electric field. Separate fits were performed at different time delays to assess the role of distinct physical processes. Only data corresponding to time delays smaller than $10s$ was used for the fits. At larger time delays, the standard deviation from the mean MSD increases considerably due to decreasing number of data points. Figure 5 shows representative plots of $MSD_{\parallel}$ and $MSD_{\perp}$ for the 70Pa, 1mA case. For all pressure-current conditions the $MSD_{\parallel}$ has a magnitude much greater than $MSD_{\perp}$ and the plots of $MSD_{\parallel}$ look almost identical to those of the combined MSDs in Figure 5. The $MSD_{\parallel}$ for all conditions has a positive concavity at all time delays, suggesting superdiffusion. The one exception is the (30 Pa, 0.35 mA) case, which has a brief negative concavity at its start. This dataset had the most noisy trajectories and the least number of data points at long time delays. Thus, the deviation in observed behavior may be due to the poorer quality of the data. The $MSD_{\perp}$ for all cases exhibits a brief negative concavity for a delay period of about two seconds, followed by a positive concavity at lager time scales. This suggests that there may be a trapping mechanism causing sub-diffusive behavior at time scales smaller than $2s$. These effects are showcased in the plots in Fig 6.
@@
Figure 7: a) $\alpha_{\parallel}$ and b) $\alpha_{\perp}$ for all pressure-current cases.
The exponent $\alpha$ extracted from $MSD_{\parallel}$ and $MSD_{\perp}$ at time delay between $4~{}s$ and $5~{}s$ is plotted for all cases in Figure 7. This time period was chosen as it has the most consistent slope fitting while not being too short of a time delay. Recall that Tsallis’ theory claims that the diffusion is a Lévy process when $q>5/3$. Using the scaling relation $MSD=\langle r^{2}\rangle\propto\tau^{\frac{2}{3-q_{p}}}$ [61] yields the criterion $\alpha>3/2$ for a Lévy process. Note that this scaling is valid only for $q_{p}$ obtained from the displacement distributions. A dashed line on Figure 7 marks the location of $\alpha=3/2$, indicating that almost all distributions for the parallel cases are a Lévy processes, while the perpendicular direction are not. Later we compare this value of $\alpha$ to that found from $q_{p}$.
@@ 5.2 Fits to Displacement Histograms
Representative histograms of the displacements are shown in Figure 9 for 70 Pa 0.7 mA case. The histograms of displacements parallel to the electric field are best approximated by a single q-Gaussian distribution, while the histograms of cross-field displacements are best approximated by a Bi-q-Gaussian distribution in most cases. These are displacements of all particles at time delay $\tau=5$. The directional difference of the distributions type implies there is anisotropic diffusion. The nonextensive $q_{p}$ parameters extracted from the displacement histograms are plotted in Figure 8.
Figure 8: Position distribution fit example from case 70 Pa 0.7 mA for displacements in a) parallel direction and b) cross field direction plotted in logarithmic scale, shifted horizontally, and zoomed in to accentuate the difference in tail behavior. The data for the position distributions is the positive squared value of $(r(\tau)-r_{0})^{2}$ instead of the velocity distributions $v_{r}$. Since the data of $(r(\tau)-r_{0})^{2}$ is already squared we fit it to a q-exponential instead of a q-Gaussian.
@@
Notice that the coefficients for the parallel direction are again larger than for the perpendicular direction. Dashed line indicates a Lévy process. The coefficients extracted from the histogram of parallel displacements suggest that two of the 70Pa cases are consistent with a Lévy process. A crossover to a Lévy process is also seen for the cross-field direction in the 70 Pa, 1 mA case if a single q-Gaussian is fitted to the histogram of cross-field displacements. However, the more accurate Bi-q-Gaussian fit to the cross-field displacement histograms suggest that the process is likely superdiffusive but not Lévy. These results are in qualitative agreement with the conclusions drown from the $\alpha$ coefficients extracted from the MSD fits at time delay $\tau=5$ shown in Figure 7. It seems that higher pressures create larger spread in the coefficients and exhibit a larger difference between the parallel and perpendicular directions suggesting a more pronounced anisotropic effect. The values for the diffusion $D\left(\frac{\mu m^{2}}{s}\right)$ calculated from Eq. 12, are shown in Figure 10. Here we can see that, in the direction parallel to the field, an increase in pressure causes diffusion to increase. However, in the cross-field direction, the interpretation is less straightforward. $D_{\perp 1}$ and $D_{\perp}$ generally decrease with pressure with some exceptions. $D_{\perp 2}$ does not show a clear trend with pressure or current.
@@
Next, we examined how the exponent $\alpha$ found directly from MSD fits compares to $\alpha$ calculated from the displacement histogram fits using the scaling in Eq. 8. Figure 11 shows the percent difference $\frac{\alpha_{MSD}-\alpha_{q_{p}}}{\alpha_{q_{p}}}$. We can see that on average the error is lower, indicated by the dashed line at 20%, for the cases where a Bi-q-Gaussian fit was used rather than a single q-Gaussian. This seems to be in agreement with findings from [28].
@@ 5.3 Velocity Distribution Plots
Next we examine Figure 12, which shows representative histograms and distribution fits for the velocity components parallel and perpendicular to the direction of the external electric field. These plots show that a Gaussian or a Maxwellian distribution (black line) does not approximate the ‘fat’ tails of the histograms, while a q-Gaussian (magenta line) and a Bi-q-Gaussian (green line) provide good fits to the $\parallel$ and the $\perp$ velocity histograms, respectively. It is again observed that the distribution features change substantially with direction, suggesting anisotropic dust diffusion caused by the external electric field. Similar to the displacement data, we observe that the velocity histograms in the cross-field direction are best approximated by a Bi-q-Gaussian fit, suggesting a superposition of two distinct diffusion processes. These trends in histogram shapes are observed for all pressure-current cases. This will be further discussed in Sec. 6.2.
@@
Kinetic temperature is typically defined as the variance of a Maxwellian distribution. Since the velocity histograms for the dust in the PK-4 experiment are best approximated by a non-Maxwellian, the notion of "temperature" here is not well defined. The range of kinetic temperatures found from the Maxwellian fits (Figure 14) is $2000K-6500K$, or about $0.2eV-0.6eV$, suggesting that the majority of the dust particles have very high kinetic energy even though the experiment is at room temperature and the pressure is low. Anomalously high dust temperature is a known phenomena that is caused by electrostatic fluctuations [67]. In Figure 15, we used Eq. 12, $T_{q}(\frac{5q-3}{2})=T_{Mq}$ (from [54]), since it provides a relation between the temperature or thermal velocity found from a q-Gaussian and to a ‘Maxwellian-like’ temperature.
@@
As shown in Figure 12 b), the perpendicular velocity histograms are best approximated by a Bi-q-Gaussian fit, where the the bulk distribution (the one with the larger variance and smaller $q_{v}$ value) more closely resembles a Gaussian and, thus, can be used to extract the thermal velocity. The second distribution in the Bi-q-Gaussian accounts for the "fat tails". Thus, in Figure 15, $T_{\perp 1}$ is extracted from a Gaussian-like sub-population distribution, while $T_{\perp 2}$ is obtained from fitting to a ’tail-halo’ distribution. It may seem unintuitive why the Gaussian sub-population would have higher temperature values, while the ’halo-tails’ have lower temperature values. This will be discussed in 6.2. Essentially we convert a variance from a distribution which has some ’tailed-ness’ to a variance which does not have any, so that we can use the same definition of temperature. We argue that this is not a different treatment of temperature in dusty plasma, but an attempt at calculating temperature more carefully with nonextensive statistics. This allows us to derive more accurate dust temperatures for the PK-4 dusty plasma with values in the range $2500K-5500K$, or about $0.2eV-0.5eV$, which is smaller than the typically calculated values of ${}10-300eV$ [67].
@@ 5.4 Domain Velocity Distribution Plots
Here we provide physical arguments why the Bi-q-Gaussian fit for the perpendicular histograms is needed. First, we consider the domain separation on Figure 3 A) which was made for the NHDS method and reconstruct the velocity histogram in each domain. Figure 16 shows the velocity distribution for domain 11 (dark blue in Figure 3 A) in both the parallel and perpendicular directions for from the 30 Pa 0.7 mA case.
@@
Table 3: The three tables show $q_{v}$ placed in the table at the corresponding domain showing a spatial map of the equilibrium.
We use the newly-founded q-Gaussian fits to calculate temperature in each domain using Eq. 12. Figure 17 a) $T_{\|}$ and b) $T_{\perp}$ show temperatures in units of Kelvin for 20 domains calculated for the 30 Pa 0.7 mA case. Each domain contained about 500 dust tracks, a minimum of 10 time frames per trajectory, which yields statistically significant amount of data for the velocity fits. The data indicates that temperatures gradients exist in the PK4 system. As seen from Eq. 16, the temperature gradient and and the gradient in potential are related through the $q_{v}$ parameter. MD simulations of ions and dust in the PK-4, with conditions closely resembling the 70 Pa 0.7 mA case, reveal gradients in the electric potential surrounding each dust particles, as shown in Figure 18 [36]. Examination of Figures 17 and 18 suggests that the dust clouds in PK-4 exhibit both temperature and electric potential gradients, in addition to q-Gaussian velocity distribution functions. This provides a qualitative argument for applying Eq. 14, which is dependent on the ratio of the thermophoretic force over the electric field force. Providing a quantitative validation for Eq. 16 is beyond the scope of the present paper and will be explored in future work.
Figure 17: a) $T_{\|}$ and b) $T_{\perp}$ calculated in each domain showing temperature gradients (30 Pa 0.7 mA).
@@ 6.1 MSD and Displacement Histograms
When analyzing the MSD plots fitting (Figure 7), it seems like both $\alpha_{\parallel}$ and $\alpha_{\perp}$ increase with pressure for dc currents 0.35 mA and 0.7 mA. The trends are less obvious at 1 mA. These fits also indicate that none of the extracted $\alpha_{\perp}$ exponents correspond to a Lévy process, while all data for $\alpha_{\parallel}$ suggest a Lévy processes. Meanwhile, the $q_{p}$ coefficients extracted from the position histogram fits (Figure 9) also suggest that $q_{p\|}$ increases with pressure (i.e., enhanced superdiffusion), in agreement with the $\alpha_{\parallel}$ trends. However, only the highest pressure-currents case (70 Pa 1mA) in Figure 9 indicates a Lévy processes (i.e., $q>5/3$). The $q_{p\perp}$ values extracted from single q-Gaussian fits show similar trends for the same currents as $\alpha_{\perp}$ but the 70Pa 1mA case suggests a Lévy process, unlike $\alpha_{\perp}$. The Bi-q-Gaussian fits yield $q_{p\perp 2}$ values that exhibit closer similarity to the $\alpha_{\perp}$ trends. We can also see from Figure 11 that the fits for $q_{p\perp 2}$ and $q_{p\perp 1}$ yield lower percent error between the predicted and measured $\alpha_{\perp}$, strengthening the argument that the Bi-q-Gaussian is a more appropriate fit. The calculated diffusion constants (Figure 10) show an overall increase in parallel diffusion with pressure, while the perpendicular diffusion seems to decrease with pressure when a single q-Gaussian fit is used. These trends are not as clear for $D_{\perp 1}$ and $D_{\perp 2}$ which were obtained from a Bi-q-Gaussian fit. Both the MSD fits and the displacement histogram fits indicate a pronounced difference in the diffusion regime when comparing the direction along the electric field versus the cross-field direction, which can be attributed to an anisotropy in the ion wake-mediated dust-dust interaction potential.